Consider the interesting curve below




which is described by the equation




cosh (sinh-¹(v²))=√1+x² dy Determine an expression for y'= (your expression will contain both x and y dx functions). Use your calculations to answer questions 1 to 4 below.




1. The derivative with regards to x of sinh ¹ (y² is




2. Using the chain rule the derivative of cosh (sinh-¹(y²)) with regards to x is




3. The derivative of √√1+x² is




dy




4. The simplified version of y'= in terms of x and y is




dx




2




3


1
Expert's answer
2021-11-24T18:34:26-0500
"\cosh(\sinh^{-1}(y^2))=\sqrt{1+x^2}"

1.


"(\sinh^{-1}(y^2))'_x=\dfrac{1}{\sqrt{1+y^4}}(2y)y'"

2.


"(\cosh(\sinh^{-1}(y^2)))'_x"

"=\sinh(\sinh^{-1}(y^2))\cdot(\sinh^{-1}(y^2))'_x"

"=y^2(\dfrac{1}{\sqrt{1+y^4}}(2y)y')=\dfrac{2y^3}{\sqrt{1+y^4}}y'"

3.


"(\sqrt{1+x^2})'_x=\dfrac{2x}{2\sqrt{1+x^2}}=\dfrac{x}{\sqrt{1+x^2}}"

4.


"\dfrac{2y^3}{\sqrt{1+y^4}}y'=\dfrac{x}{\sqrt{1+x^2}}"

"y'=\dfrac{x\sqrt{1+y^4}}{2y^3\sqrt{1+x^2}}"